HI,
You see I have homework telling me to find all the different possible perimeters for 36 and I would like some help.
Thanks.

Answer by Mathtut(3670)

since perimeter = 2 times the length +2 times the width
:
2L+2W=P......2L+2W=36....there is an infinite amount of solutions....if you looking for some examples just plug in any number for either variable and solve for the other.
:
such as L=1....2(1)+2W=36....2W=34...W=17
or
L=4 2(4)+2W=36...2W=28...W=14
this can go on forever...